Given the following function of the monopolist: average cost= 140+9Q+5Q2
Average revenue= 170 + 16Q + Q2
Calculate total cost, Total Revenue,marginal cost, marginal revenue, equilibrium output, profit or loss
1
Expert's answer
2021-08-02T11:05:36-0400

Total cost is:

"TC = AC\u00d7Q = 140Q + 9Q^2 + 5Q^3."

Total Revenue is:

"TR = AR\u00d7Q = 170Q + 16Q^2 + Q^3."

Marginal cost is:

"MC = TC'(Q) = 140 + 18Q + 15Q^2."

Marginal revenue is:

"MR = TR'(Q) = 170 + 32Q + 3Q^2."

Equilibrium output is:

MR = MC,

"170 + 32Q + 3Q^2 = 140 + 18Q + 15Q^2,"

"6Q^2 - 7Q - 15 = 0,"

Q = (7 + 20.22)/12 = 2.27 units.

Profit is:

"TP = TR - TC = 170\u00d72.27 + 16\u00d72.27^2 + 2.27^3 - (140\u00d72.27 + 9\u00d72.27^2 + 5\u00d72.27^3) = 57.38."